OpenStudy (anonymous):

What is the solution of the following system? 3x+3y= −9x−9y=-30

OpenStudy (gudden):

Is it 3x +3y = - 9x- 9y = 30 ??

OpenStudy (gudden):

@mamatodylan .... you there??

OpenStudy (anonymous):

3x+3y=10 -9x-9y=-30

TheSmartOne (thesmartone):

basically the two equations are the same thing because if you divide the second equation by-3 you get teh first one.